OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? ] y = 3/2x y = -1/2x + 4 ( , )

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

solve \[\frac{3}{2}x=-\frac{1}{2}x+4\] for \(x\)

OpenStudy (misty1212):

easier if you multiply both sides by 2 and start with \[3x=-x+8\]

OpenStudy (misty1212):

add x to both sides, then divide both sides by 4

OpenStudy (anonymous):

Put y=3/2x in second eqaution , then solve for x , after that u will get x . Then put value of x in any above equation , u will get value of y ... that will be ur solution (x , y)

OpenStudy (anonymous):

3 divided by 4 = 0.75 and 8 divided by 4 = 4 so is it 0.75 for x and 4 for y?

OpenStudy (anonymous):

no please check again ... first u have 3/2x = -1/2x +4 Then just multiply both sides by 2 , u will get 3x = -x + 8 okay ?

OpenStudy (anonymous):

ummm XD

OpenStudy (anonymous):

would it be X = 3 and 8 = y?

OpenStudy (anonymous):

wait , we have got 3x = -x + 8 now can u move -x to left hand side ,,,? ( hint just add x to both sides)

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

So what do u get ?

OpenStudy (anonymous):

4 = 4?

OpenStudy (anonymous):

see the step here we have got 3x = -x + 8 now to move -x to left hand side we just add x to both sides so we get 3x + x = -x +x +8 => 4x = 8 okay?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so x = 2?

OpenStudy (anonymous):

Correct !!

OpenStudy (anonymous):

Now just put this value x=2 in any above equation u will get y ..

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

1.5 * 2 = 3

OpenStudy (anonymous):

Very good !!!

OpenStudy (anonymous):

0.5 * 2 = 1 + 4 = 5 so umm

OpenStudy (anonymous):

oh its -1/2 XD

OpenStudy (anonymous):

So we have got (2 , 3)

OpenStudy (anonymous):

ye its 3 :D thanks for the help

OpenStudy (anonymous):

U r Welcome :)